15=30 I don't know how to approch this problem.
New Toy Company is producing a raing car set, complete with plastic track. The track is a solid surface and has a straight edge with circular ends. See the following sketch for dimensions. What is the distance around the outside of track? (Use 3.14 to approximate Pi)
r= 7 and the striaght edge 15 inches.

Answer by checkley77(12844)

Straight runs=2*15=30 inches.
The 2 semi-circular ends make a complete circle with a radius of 7 inches.
c=2*pi*7
c=2*3.14*7
c=6.24*7
c=43.68 inches for the 2 end semi circles.
43.68+30=73.68 total inches around the oval track.